OpenStudy (anonymous):

A sledgehammer weighs 18 pounds on Earth and 4 pounds on the planet Namar. Based on the numbers, an 80–pound earthling would weigh about _____ pounds on Namar.

OpenStudy (texaschic101):

18/4 = 80/x (18 earth lbs to 4 namar lbs = 80 earth lbs to x namar lbs now cross multiply (18)(x) = (4)(80) now do the math and solve for x.

OpenStudy (anonymous):

hey i ryd to but it would give me wrong numbers

OpenStudy (texaschic101):

18x = 320 x = 320/18 x = 17.77 rounds to 18 lbs

OpenStudy (anonymous):

A sledgehammer weighs 18 pounds on Earth and 4 pounds on the planet Namar. Based on the numbers given, your weight on Namar would be about _____ times your weight on Earth.

OpenStudy (anonymous):

360